OpenStudy (anonymous):

x-2= square root2x-5 solve equation???

Parth (parthkohli):

Can you find the square of both sides?

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x - 2)^2 = \sqrt{2x - 5}^2}\)

Parth (parthkohli):

Remember: \((a - b)^2 = a^2 - 2ab + b^2\)

OpenStudy (anonymous):

(a-b)^2 = a^2 - 2ab + b^2

OpenStudy (anonymous):

2x-5=x^2-4x+4

Parth (parthkohli):

That's right. Now add 5 to both sides, then subtract 2x from both.

Parth (parthkohli):

You'd supposedly have a quadratic equation after doing all that.

OpenStudy (theviper):

Square both sides :)

OpenStudy (anonymous):

x^2-6x +9

Parth (parthkohli):

That's right! You may use factoring, as that is a perfect square.

Parth (parthkohli):

Factoring/completing the square.

OpenStudy (anonymous):

answer would be (-3,3)

Parth (parthkohli):

Nope, there's only one root here with a multiplicity of 2(remember how we have an exponent of 2?)

Parth (parthkohli):

You may express \(x^2 - 6x + 9\) as a perfect square. How'd you do that?

OpenStudy (anonymous):

(x-3)(x+3)

Parth (parthkohli):

Nope. That makes \(x^2 - 9\) when you multiply 'em out.

Parth (parthkohli):

And that's not a perfect square..

OpenStudy (anonymous):

your right.

OpenStudy (anonymous):

(x-3)(x-3)

OpenStudy (anonymous):

so x=3

Parth (parthkohli):

Yep, so what's the solution?

OpenStudy (anonymous):

x=3

Parth (parthkohli):

That's correct!